Smith Company produces and sells one product for $40 per unit. The company has no

beginning inventories. Its variable manufacturing cost per unit is $18 and the variable

selling and administrative expense per unit is $4. The fixed manufacturing overhead and

fixed selling and administrative expense total $80,000 and $20,000, respectively. If

Smith Company produces 8,000 units and sells 7,500 units during the year, then its net

operating income under absorption and variable costing
Answer #1
Absorption costing:
Fixed manufacturing overhead per unit 10 =80000/8000
Sales revenue 300000 =7500*40
Less: Cost of goods sold 210000 =7500*(18+10)
Gross Margin 90000
Less: Selling and administrative expense 50000 =20000+(7500*4)
Net operating income under absorption costing 40000
Variable costing:
Sales revenue 300000 =7500*40
Less: Variable Cost of goods sold 135000 =7500*18
Manufacturing margin 165000
Less: Variable selling and administrative expense 30000 =7500*4
Contribution Margin 135000
Less: Fixed manufacturing overhead 80000
Less: Fixed Selling and administrative expense 20000
Net operating income under variable costing 35000